# Training Robust Graph Neural Networks by Modeling Noise Dependencies

### v1

@@ 4.2.2 Modeling Generative Decoder
Modeling $p_{\theta_{1}}(A|X,\epsilon,Z_{A})$. The probability $p(A|X,\epsilon,Z_{A})$ means the likelihood of how well the noisy edge $A$ is reconstructed from the latent graph structure $Z_{A}$ along with $\epsilon$ and $X$. Hence, we aim to minimize $-\mathbb{E}_{Z_{A}\sim q_{\phi_{1}}}\mathbb{E}_{\epsilon\sim q_{\phi_{2}}}\left[\log(p_{\theta_{1}}(A|X,\epsilon,Z_{A}))\right]$ to discover the latent graph structure $Z_{A}$ from which the noisy edge $A$ is reconstructed given noise sources, $X$ and $\epsilon$. We implement it as an edge reconstruction loss forcing the estimated latent structure $\hat{\mathbf{A}}$ to assign greater weights to clean edges and reduce the influence of noisy edges, which is defined as ${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\mathcal{L}_{\text{rec-edge}}}$: which is defined as ${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\mathcal{L}_{\text{rec-edge}}}$:
$\small{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\mathcal{L}_{\text{rec-edge}}}=\frac{N}{|\mathcal{E}|+|\mathcal{E^{-}}|}\left(\sum_{(i,j)\in\mathcal{E}}(\hat{p}^{reg}_{ij}-\hat{p}^{el}_{ij})^{2}+\sum_{(i,j)\in\mathcal{E}^{-}}(\hat{p}_{ij}-0)^{2}\right),$ (2)
@@
More precisely, the regularized prediction $\hat{p}^{reg}_{ij}$ is defined as: $\hat{p}^{reg}_{ij}=\theta_{1}\hat{p}_{ij}+(1-\theta_{1})s(\mathbf{X}_{i},\mathbf{X}_{j})$. The main idea is to penalize $\hat{p}_{ij}$ when $s(\mathbf{X}_{i},\mathbf{X}_{j})$ is high, as the edge between $v_{i}$ and $v_{j}$ is potentially noisy due to the influence of noisy $X$. To regularize labels, we adopt label smoothing approach by $\hat{p}_{ij}^{el}\in[0.9,1]$, enhancing the robustness in the presence of noisy supervision. When an edge is regarded as noisy (i.e., with a low $\hat{p}_{ij}^{el}$), its label is close to 0.9333The value 0.9 is selected following (Szegedy et al., 2016)., while an edge considered clean (i.e., with a high $\hat{p}_{ij}^{el}$) has a label close to 1.
Modeling $p_{\theta_{2}}(X|\epsilon,Z_{Y}))$. The term $p(X|\epsilon,Z_{Y})$ indicates how well the noisy node feature $X$ is reconstructed from the latent clean label $Z_{Y}$ along with $\epsilon$. Hence, we aim to minimize $-\mathbb{E}_{\epsilon\sim q_{\phi_{2}}}\mathbb{E}_{Z_{Y}\sim q_{\phi_{3}}}\left[\log(p_{\theta_{2}}(X|\epsilon,Z_{Y}))\right]$. To do so, the decoder needs to rely on the information contained in $Z_{Y}$, which essentially encourages the value of $Z_{Y}$ to be meaningful for the prediction process, i.e., generating $X$. It is implemented as a feature reconstruction loss ${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\mathcal{L}_{\text{rec-feat}}}$, where the decoder $\theta_{2}$ is composed of an MLP that takes $\epsilon_{X}$ and $Z_{Y}$ as inputs and reconstructs node features. Note that the reparametrization trick (Kingma & Welling, 2013) is used for sampling $\epsilon_{X}$ that follows the standard normal distribution.
Modeling $p_{\theta_{3}}(Y|X,A,Z_{Y})$. The term $p(Y|X,A,Z_{Y})$ means the transition relationship from the latent clean label $Z_{Y}$ to the noisy label $Y$ of an instance, i.e., how the label noise was generated (Yao et al., 2021). For this reason, maximizing $\log(p_{\theta_{3}}(Y|X,A,Z_{Y}))$ would let us discover the latent true label $Z_{Y}$ from which the noisy label $Y$ is generated given an instance, i.e., $X$ and $A$. Hence, we aim to maximize the log likelihood, which is implemented as minimizing a node classification loss $\mathcal{L}_{\text{cls-dec}}$. Specifically, the decoder $\theta_{3}$ is composed of a GCN classifier: $\hat{\mathbf{Y}}_{\text{dec}}=\text{GCN}_{\theta_{3}}(\mathbf{X},\mathbf{A},\mathbf{\hat{Y}})\in\mathbb{R}^{N\times C}$. Note that such a learning objective is equivalent to minimizing $-\mathbb{E}_{Z_{Y}\sim q_{\phi_{3}}}\left[\log(p_{\theta_{3}}(Y|X,A,Z_{Y}))\right]$ in Eqn. 1.
@@ 4.2.3 Model Training
$\small\mathcal{L}_{\text{final}}=\mathcal{L}_{\text{cls-enc}}+\lambda_{1}\mathcal{L}_{\text{rec-edge}}+\lambda_{2}\mathcal{L}_{\text{hom}}+\lambda_{3}(\mathcal{L}_{\text{rec-feat}}+\mathcal{L}_{\text{cls-dec}}+\mathcal{L}_{\text{p}}),$ (3)
@@ 5 Experiments
Datasets. We evaluate Da-Gnn and baselines on five commonly used benchmark datasets and two newly introduced datasets, Auto and Garden, which are generated upon Amazon review data (He & McAuley, 2016; McAuley et al., 2015) to mimic DANG on e-commerce systems (Refer to Appendix E.2.2 for details). The details of the datasets are given in Appendix E.1.
Experimental Details. We evaluated Da-Gnn in both node classification and link prediction tasks, comparing it with noise-robust GNNs and generative GNN methods. For a thorough evaluation, we create synthetic and real-world DANG benchmark datasets, with details in Appendix E.2. We also account for other noise scenarios, commonly considered in this research field, following (Li et al., 2022a; Liu et al., 2021; Qian et al., 2023). Further details about the baselines, evaluation protocol, and implementation details can be found in Appendix E.3, E.4, and E.5, respectively.
@@
5.1 Main Results
2) Da-Gnn also shows comparable or better performance than baselines under other noise scenarios, commonly considered in this research field. Specifically, we evaluate the robustness of Da-Gnn under commonly utilized node feature noise (Liu et al., 2021), structure noise (Li et al., 2022a), and node label noise scenarios (Dai et al., 2021) on Cora dataset [^4] Additional results on other datasets are outlined in Appendix F.1. In Fig 4, we observe Da-Gnn shows consistent superiority or competitive performance compared to existing robust GNNs. We attribute the robustness of Da-Gnn under the noise in node features to the graph structure learning module that accurately infers the latent graph structure $Z_{A}$. The utilization of abundant local neighborhoods acquired through the inference of $Z_{A}$ enables effective smoothing for nodes with noisy features, leveraging the information within these neighborhoods. We attribute the effectiveness of Da-Gnn under the noise in graph structures to inferring the robust latent clean graph structure. In other words, the inference of the latent clean graph structure $Z_{A}$ assigns greater weights to latent clean edges and lower weights to observed noisy edges by employing regularizations on both the edge predictions and labels, thereby mitigating structural noise. For the noise in node labels, we argue that the effectiveness of Da-Gnn stems from the accurate inference of the latent clean structure. Specifically, the inferred latent node label $Z_{Y}$ is regularized using the inferred latent structure $Z_{A}$ to meet the homophily assumption (i.e., $\mathcal{L}_{\text{hom}}$). Leveraging the clean neighbor structure, this regularization technique has been demonstrated to effectively address noisy labels (Iscen et al., 2022).
@@ 5.1 Main Results
Figure 5: Node classification under extreme noise scenario.
Figure 6: Graphical models of DGPs derived from DANG.
Table 3: Ablation studies regarding various DGPs in Fig 7. In Case 3, the causal relationship $Y\leftarrow(X,A)$ is removed from the DGP of DANG (See Fig 2(b)). In Case 2, $A\leftarrow X$ is additionally removed. In Case 1, $A\leftarrow\epsilon$ is additionally removed, which is equivalent to the DGP of IFN (See Fig 2(a)).
Dataset Setting (a) Case 1 (b) Case 2 (c) Case 3 Da-Gnn
@@
5.2 Ablation Studies
@@
1) Removing the causal relationships $Y\leftarrow(X,A)$, i.e., Fig 7(c), simplifies $-\mathbb{E}_{Z_{Y}\sim q_{\phi_{3}}}\left[\log(p_{\theta_{3}}(Y|X,A,Z_{Y}))\right]$ to $-\mathbb{E}_{Z_{Y}\sim q_{\phi_{3}}}\left[\log(p_{\theta_{3}}(Y|Z_{Y}))\right]$. This simplification hinders the accurate modeling of the label transition relationship from $Z_{Y}$ to the noisy label $Y$, resulting in a degradation of model performance under DANG.
2) Additionally, when excluding the causal relationship $A\leftarrow X$, i.e., Fig 7(b), the inference of $Z_{A}$ and $Z_{Y}$ is simplified as follows: $q_{\phi_{1}}(Z_{A}|X,A)$ to $q_{\phi_{1}}(Z_{A}|A)$ and $q_{\phi_{3}}(Z_{Y}|X,A)$ to $q_{\phi_{3}}(Z_{Y}|X)$. Furthermore, the loss term $-\mathbb{E}_{Z_{A}\sim q_{\phi_{1}}}\mathbb{E}_{\epsilon\sim q_{\phi_{2}}}\left[\log(p_{\theta_{1}}(A|X,\epsilon,Z_{A}))\right]$ is also simplified to $-\mathbb{E}_{Z_{A}\sim q_{\phi_{1}}}\mathbb{E}_{\epsilon\sim q_{\phi_{2}}}\left[\log(p_{\theta_{1}}(A|\epsilon,Z_{A}))\right]$. These simplifications significantly hinder the accurate inference of $Z_{A}$ and $Z_{Y}$, resulting in a notable performance degradation.
3) Eliminating the causal relationship $A\leftarrow\epsilon$, as in Fig 7(a), simplifies $-\mathbb{E}_{Z_{A}\sim q_{\phi_{1}}}\mathbb{E}_{\epsilon\sim q_{\phi_{2}}}\left[\log(p_{\theta_{1}}(A|\epsilon,Z_{A}))\right]$ to $-\mathbb{E}_{Z_{A}\sim q_{\phi_{1}}}\mathbb{E}_{\epsilon\sim q_{\phi_{2}}}\left[\log(p_{\theta_{1}}(A|Z_{A}))\right]$. This simplification hinders the robustness of the inferred $Z_{A}$, since the simplified loss excludes label regularization from the model training process, ultimately resulting in performance degradation.
5.3 Qualitative Analysis
5.4 Complexity Analysis
@@ 5.5 Sensitivity Analysis
We analyze the sensitivity of hyperparameters $\lambda_{1}$, $\lambda_{2}$, $\theta$, and $\gamma$. Our observations indicate that while each hyperparameter reaches an optimal point, the overall performance of Da-Gnn consistently surpasses that of the state-of-the-art baseline and demonstrates robustness to variations in hyperparameter values. For a more comprehensive discussion and detailed results, please see Appendix F.4.
@@ 7 Limitations and Future Works
While the proposed DANG and Da-Gnn demonstrate broader applicability compared to existing works, they do not perfectly cover all possible noise scenarios. One potential direction to enhance their practicality is to incorporate the causal relationship $X\leftarrow A$, which suggests that graph structure noise can inevitably lead to node feature noise—an occurrence that may manifest in certain real-world scenarios. By accounting for this relationship, a broader range of noise scenarios could be addressed, further improving practical applicability. A detailed discussion on this topic is provided in Appendix C.2.
8 Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Derivation Details of Evidence Lower BOund (ELBO)
For simplicity, we introduce two additional assumptions. First, when the node features $X$ and observed graph structure $A$ are given, latent clean graph structure $Z_{A}$ is conditionally independent from the noise-incurring variable $\epsilon$, i.e., $q_{\phi_{1}}(Z_{A}|X,A,\epsilon)=q_{\phi_{1}}(Z_{A}|X,A)$. Second, when $X$ and $A$ are given, latent clean labels $Z_{Y}$ is conditionally independent from the observed node labels $Y$, i.e., $q_{\phi_{3}}(Z_{Y}|X,A,Y)=q_{\phi_{3}}(Z_{Y}|X,A)$. This approximation, known as the mean-field method, is a prevalent technique utilized in variational inference-based methods (Ma et al., 2019; Lao et al., 2022). As a result, we can simplify Eqn. 6 as follows:
@@
To jointly optimize the parameter $\phi$ and $\theta$, we adopt the variational inference framework (Blei et al., 2017; Yao et al., 2021) to optimize the Evidence Lower-BOund (ELBO) of the marginal likelihood for observed data, rather than optimizing the marginal likelihood directly. Specifically, we derive the ELBO for the observed data log-likelihood $P(X,A,Y)$. First, we factorize the joint distribution $P(X,A,Y,\epsilon,Z_{A},Z_{Y})$ based on the graphical model in Fig. 2(b) in the main paper:
@@
Recall that the conditional distribution $q_{\phi}(\epsilon,Z_{A},Z_{Y}|X,A,Y)$ is factorized as in Eqn. 7. Now, we derive the ELBO for the observed data log-likelihood $P(X,A,Y)$:
[…]
@@
.$\biggl{.}\quad\quad\quad\quad\quad\quad\quad\quad\quad+\log(p_{\theta_{2}}(X|\epsilon,Z_{Y}))+\log(p_{\theta_{3}}(Y|X,A,Z_{Y}))\biggr{]}$
@@
The last equation of Eq. 10 can be more simplified. We present the simplified results in Eqn. 11, 12, 13, and 14, where we abuse the notation $\mathbb{E}_{Z_{A}\sim q_{\phi_{1}}(Z_{A}|X,A)}$, $\mathbb{E}_{\epsilon\sim q_{\phi_{2}}(\epsilon|X,A,Z_{Y})}$, and $\mathbb{E}_{Z_{Y}\sim q_{\phi_{3}}(Z_{Y}|X,A)}$ as $q_{\phi_{1}}$, $q_{\phi_{2}}$, and $q_{\phi_{3}}$, respectively:
@@
We combine Eqn. 11, 12, 13, and 14 to get the negative ELBO, i.e., $\mathcal{L}_{\text{ELBO}}$:

### v2

@@ 4.2.2 Modeling Generative Decoder
Modeling $p_{\theta_{1}}(A|X,\epsilon,Z_{A})$. The probability $p(A|X,\epsilon,Z_{A})$ means the likelihood of how well the noisy edge $A$ is reconstructed from the latent graph structure $Z_{A}$ along with $\epsilon$ and $X$. Hence, we aim to minimize $\mathbb{E}_{Z_{A}\sim q_{\phi_{1}}}\mathbb{E}_{\epsilon\sim q_{\phi_{2}}}\left[\log(p_{\theta_{1}}(A|X,\epsilon,Z_{A}))\right]$ to discover the latent graph structure $Z_{A}$ from which the noisy edge $A$ is reconstructed given noise sources, $X$ and $\epsilon$. We implement it as an edge reconstruction loss forcing the estimated latent structure $\hat{\mathbf{A}}$ to assign greater weights to clean edges and reduce the influence of noisy edges, which is defined as $\color[rgb]{.75,0,.25}\mathcal{L}_{\text{rec-edge}}$}: which is defined as $\color[rgb]{.75,0,.25}\mathcal{L}_{\text{rec-edge}}$}:
${\color[rgb]{.75,0,.25}\mathcal{L}_{\text{rec-edge}}}=\frac{N}{|\mathcal{E}|+|\mathcal{E^{-}}|}\left(\sum_{(i,j)\in\mathcal{E}}(\hat{p}^{reg}_{ij}-\hat{p}^{el}_{ij})^{2}+\sum_{(i,j)\in\mathcal{E}^{-}}(\hat{p}_{ij}-0)^{2}\right),$ (3)
@@
More precisely, the regularized prediction $\hat{p}^{reg}_{ij}$ is defined as: $\hat{p}^{reg}_{ij}=\theta_{1}\hat{p}_{ij}+(1-\theta_{1})s(\mathbf{X}_{i},\mathbf{X}_{j})$. The main idea is to penalize $\hat{p}_{ij}$ when $s(\mathbf{X}_{i},\mathbf{X}_{j})$ is high, as the edge between $v_{i}$ and $v_{j}$ is potentially noisy due to the influence of noisy $X$. To regularize labels, we adopt label smoothing approach by $\hat{p}_{ij}^{el}\in[0.9,1]$, enhancing the robustness in the presence of noisy supervision. When an edge is regarded as noisy (i.e., with a low $\hat{p}_{ij}^{el}$), its label is close to 0.9333The value 0.9 is selected following inception ., while an edge considered clean (i.e., with a high $\hat{p}_{ij}^{el}$) has a label close to 1.
Modeling $p_{\theta_{2}}(X|\epsilon,Z_{Y}))$. The term $p(X|\epsilon,Z_{Y})$ indicates how well the noisy node feature $X$ is reconstructed from the latent clean label $Z_{Y}$ along with $\epsilon$. Hence, we aim to minimize $\mathbb{E}_{\epsilon\sim q_{\phi_{2}}}\mathbb{E}_{Z_{Y}\sim q_{\phi_{3}}}\left[\log(p_{\theta_{2}}(X|\epsilon,Z_{Y}))\right]$. To do so, the decoder needs to rely on the information contained in $Z_{Y}$, which essentially encourages the value of $Z_{Y}$ to be meaningful for the prediction process, i.e., generating $X$. It is implemented as a feature reconstruction loss $\color[rgb]{.75,0,.25}\mathcal{L}_{\text{rec-feat}}$}, where the decoder $\theta_{2}$ is composed of an MLP that takes $\epsilon_{X}$ and $Z_{Y}$ as inputs and reconstructs node features. Note that the reparametrization trick kingma2013auto is used for sampling $\epsilon_{X}$ that follows the standard normal distribution.
Modeling $p_{\theta_{3}}(Y|X,A,Z_{Y})$. The term $p(Y|X,A,Z_{Y})$ means the transition relationship from the latent clean label $Z_{Y}$ to the noisy label $Y$ of an instance, i.e., how the label noise was generated yao2021instance . For this reason, maximizing $\log(p_{\theta_{3}}(Y|X,A,Z_{Y}))$ would let us discover the latent true label $Z_{Y}$ from which the noisy label $Y$ is generated given an instance, i.e., $X$ and $A$. Hence, we aim to maximize the log likelihood, which is implemented as minimizing a node classification loss $\mathcal{L}_{\text{cls-dec}}$. Specifically, the decoder $\theta_{3}$ is composed of a GCN classifier: $\hat{\mathbf{Y}}_{\text{dec}}=\text{GCN}_{\theta_{3}}(\mathbf{X},\mathbf{A},\mathbf{\hat{Y}})\in\mathbb{R}^{N\times C}$. Note that such a learning objective is equivalent to minimizing $\mathbb{E}_{Z_{Y}\sim q_{\phi_{3}}}\left[\log(p_{\theta_{3}}(Y|X,A,Z_{Y}))\right]$ in Eqn. 1.
@@ 4.2.3 Model Training
$\mathcal{L}_{\text{final}}=\mathcal{L}_{\text{cls-enc}}+\lambda_{1}\mathcal{L}_{\text{rec-edge}}+\lambda_{2}\mathcal{L}_{\text{hom}}+\lambda_{3}(\mathcal{L}_{\text{rec-feat}}+\mathcal{L}_{\text{cls-dec}}+\mathcal{L}_{\text{p}}),$ (4)
@@ 5 Experiments
Datasets. We evaluate DA-GNN and baselines on five commonly used benchmark datasets and two newly introduced datasets, Auto and Garden, which are generated upon Amazon review data amazondata1 ; amazondata2 to mimic DANG on e-commerce systems (Refer to Appendix E.2.2 for details). The details of the datasets are given in Appendix E.1.
Experimental Details. We evaluated DA-GNN in both node classification and link prediction tasks, comparing it with noise-robust GNNs and generative GNN methods. For a thorough evaluation, we create synthetic and real-world DANG benchmark datasets, with details in Appendix E.2. We also account for other noise scenarios, commonly considered in this research field, following stable ; airgnn ; rtgnn . Further details about the baselines, evaluation protocol, and implementation details can be found in Appendix E.3, E.4, and E.5, respectively.
@@ 5.1 Main Results
2) DA-GNN also shows comparable or better performance than baselines under other noise scenarios, commonly considered in this research field. Specifically, we evaluate the robustness of DA-GNN under commonly utilized node feature noise airgnn , structure noise stable , and node label noise scenarios nrgnn on Cora dataset [^4] Additional results on other datasets are outlined in Fig 12, 12, and 12 in Appendix.. In Fig 4, we observe DA-GNN shows consistent superiority or competitive performance compared to existing robust GNNs. We attribute the robustness of DA-GNN under the noise in node features to the graph structure learning module that accurately infers the latent graph structure $Z_{A}$. The utilization of abundant local neighborhoods acquired through the inference of $Z_{A}$ enables effective smoothing for nodes with noisy features, leveraging the information within these neighborhoods. We attribute the effectiveness of DA-GNN under the noise in graph structures to inferring the robust latent clean graph structure. In other words, the inference of the latent clean graph structure $Z_{A}$ assigns greater weights to latent clean edges and lower weights to observed noisy edges by employing regularizations on both the edge predictions and labels, thereby mitigating structural noise. For the noise in node labels, we argue that the effectiveness of DA-GNN stems from the accurate inference of the latent clean structure. Specifically, the inferred latent node label $Z_{Y}$ is regularized using the inferred latent structure $Z_{A}$ to meet the homophily assumption (i.e., $\mathcal{L}_{\text{hom}}$). Leveraging the clean neighbor structure, this regularization technique has been demonstrated to effectively address noisy labels iscen2022consistency .
Figure 5: Node classification under extreme noise scenario.
@@
5.2 Ablation Studies on DA-GNN
Table 3: Ablation studies of various DGPs from Fig 6. Case 3 removes $Y\leftarrow(X,A)$; Case 2 additionally removes $A\leftarrow X$; Case 1 additionally removes $A\leftarrow\epsilon$, equivalent to IFN (Fig 2).
Dataset Setting (a) Case 1 (b) Case 2 (c) Case 3 Proposed
@@ 5.2 Ablation Studies on DA-GNN
Figure 6: Graphical models of DGPs derived from DANG.
@@
1) Removing $Y\leftarrow(X,A)$, i.e., Fig 6(c), simplifies $\mathbb{E}_{Z_{Y}\sim q_{\phi_{3}}}\left[\log(p_{\theta_{3}}(Y|X,A,Z_{Y}))\right]$ to $\mathbb{E}_{Z_{Y}\sim q_{\phi_{3}}}\left[\log(p_{\theta_{3}}(Y|Z_{Y}))\right]$. This simplification hinders the accurate modeling of the label transition relationship from $Z_{Y}$ to the noisy label $Y$, resulting in a degradation of model performance under DANG.
2) Additionally, when excluding $A\leftarrow X$, i.e., Fig 6(b), the inference of $Z_{A}$ and $Z_{Y}$ is simplified as follows: $q_{\phi_{1}}(Z_{A}|X,A)$ to $q_{\phi_{1}}(Z_{A}|A)$ and $q_{\phi_{3}}(Z_{Y}|X,A)$ to $q_{\phi_{3}}(Z_{Y}|X)$. Furthermore, the loss term $\mathbb{E}_{Z_{A}\sim q_{\phi_{1}}}\mathbb{E}_{\epsilon\sim q_{\phi_{2}}}\left[\log(p_{\theta_{1}}(A|X,\epsilon,Z_{A}))\right]$ is also simplified to $\mathbb{E}_{Z_{A}\sim q_{\phi_{1}}}\mathbb{E}_{\epsilon\sim q_{\phi_{2}}}\left[\log(p_{\theta_{1}}(A|\epsilon,Z_{A}))\right]$. These simplifications significantly hinder the accurate inference of $Z_{A}$ and $Z_{Y}$, resulting in a notable performance degradation.
3) Eliminating $A\leftarrow\epsilon$, as in Fig 6(a), simplifies $\mathbb{E}_{Z_{A}\sim q_{\phi_{1}}}\mathbb{E}_{\epsilon\sim q_{\phi_{2}}}\left[\log(p_{\theta_{1}}(A|\epsilon,Z_{A}))\right]$ to $\mathbb{E}_{Z_{A}\sim q_{\phi_{1}}}\mathbb{E}_{\epsilon\sim q_{\phi_{2}}}\left[\log(p_{\theta_{1}}(A|Z_{A}))\right]$. This simplification hinders the robustness of the inferred $Z_{A}$, since the simplified loss excludes label regularization from the model training process, ultimately resulting in performance degradation.
5.3 Complexity Analysis on DA-GNN
@@ 5.4 Sensitivity Analysis
We analyze the sensitivity of our proposed method DA-GNN in terms of its hyperparameters $\lambda_{1}$, $\lambda_{2}$, $k$, $\theta$, and $\gamma$. Our observations indicate that DA-GNN consistently exhibit best performance regardless of their values. Among these, $k$ plays a critical role and requires some tuning. But, as the search space is relatively small, we consider this acceptable. For a more comprehensive discussion and detailed results, please see Appendix F.2.
5.5 Robustness Evaluation under Variants of DANG
We analyze the robustness of DA-GNN across different variants of DANG by varying the hyperparameter settings used in dataset generation. Specifically, in the generation process of our synthetic DANG, we have three variables: 1) the overall noise rate, 2) the amount of noise dependency ($X\rightarrow A$, ($X\rightarrow Y$, ($A\rightarrow Y$), and 3) the amount of independent structure noise ($\epsilon\rightarrow A$). For the generation process of our real-world DANG, we have 1) the number of fraudsters (i.e., nodes with noisy features) and 2) the activeness of fraudsters (i.e., the amount of structure noise they introduce). As a result, label noise also increases accordingly, in proportion to the amount of generated feature and structure noise.
Detailed results in Appendix F.3 show that DA-GNN consistently outperforms all baselines across varying levels of both synthetic and real-world DANG, underscoring its robustness and practical applicability under diverse noise conditions.
5.6 Qualitative Analysis on DA-GNN
We conduct qualitative analyses to verify how well DA-GNN infers the latent variables $\epsilon_{A}$ and $Z_{A}$. For a detailed setting and results, please refer to Appendix F.4.
@@ 7 Limitations and Future Works
Despite broader applicability of the DANG and DA-GNN, they do not perfectly cover all possible noise scenarios. One direction to enhance their practicality is to incorporate $X\leftarrow A$, suggesting graph structure noise can inevitably lead to node feature noise. By doing so, a broader range of noise scenarios could be addressed, further improving practical applicability. A detailed discussion on this topic is provided in Appendix C.3.
Acknowledgements
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2025-02304967), IITP grant funded by the Korea government(MSIT) (RS-2022-II220157), National Research Foundation of Korea(NRF) funded by Ministry of Science and ICT (RS-2022-NR068758).
@@ Appendix A Derivation Details of Evidence Lower BOund (ELBO)
For simplicity, we introduce two additional assumptions. First, when the node features $X$ and observed graph structure $A$ are given, latent clean graph structure $Z_{A}$ is conditionally independent from the noise-incurring variable $\epsilon$, i.e., $q_{\phi_{1}}(Z_{A}|X,A,\epsilon)=q_{\phi_{1}}(Z_{A}|X,A)$. Second, when $X$ and $A$ are given, latent clean labels $Z_{Y}$ is conditionally independent from the observed node labels $Y$, i.e., $q_{\phi_{3}}(Z_{Y}|X,A,Y)=q_{\phi_{3}}(Z_{Y}|X,A)$. This approximation, known as the mean-field method, is a prevalent technique utilized in variational inference-based methods ma2019flexible ; lao2022variational . As a result, we can simplify Eqn. 7 as follows:
@@
To jointly optimize the parameter $\phi$ and $\theta$, we adopt the variational inference framework vi_review ; yao2021instance to optimize the Evidence Lower-BOund (ELBO) of the marginal likelihood for observed data, rather than optimizing the marginal likelihood directly. Specifically, we derive the ELBO for the observed data log-likelihood $P(X,A,Y)$. First, we factorize the joint distribution $P(X,A,Y,\epsilon,Z_{A},Z_{Y})$ based on the graphical model in Fig. 2(b) in the main paper:
@@
Recall that the conditional distribution $q_{\phi}(\epsilon,Z_{A},Z_{Y}|X,A,Y)$ is factorized as in Eqn. 8. Now, we derive the ELBO for the observed data log-likelihood $P(X,A,Y)$:
$\log p_{\theta}(X,A,Y)=\log\int_{\epsilon}\int_{Z_{A}}\int_{Z_{Y}}p_{\theta}(X,A,Y,\epsilon,Z_{A},Z_{Y})d\epsilon dZ_{A}dZ_{Y}$
@@
$\biggl.\quad\quad\quad\quad\quad\quad\quad\quad\quad+\log(p_{\theta_{2}}(X|\epsilon,Z_{Y}))+\log(p_{\theta_{3}}(Y|X,A,Z_{Y}))\biggr]$
@@
The last equation of Eq. 11 can be more simplified. We present the simplified results in Eqn. 12, 13, 14, and 15, where we abuse the notation $\mathbb{E}_{Z_{A}\sim q_{\phi_{1}}(Z_{A}|X,A)}$, $\mathbb{E}_{\epsilon\sim q_{\phi_{2}}(\epsilon|X,A,Z_{Y})}$, and $\mathbb{E}_{Z_{Y}\sim q_{\phi_{3}}(Z_{Y}|X,A)}$ as $q_{\phi_{1}}$, $q_{\phi_{2}}$, and $q_{\phi_{3}}$, respectively:
@@
We combine Eqn. 12, 13, 14, and 15 to get the negative ELBO, i.e., $\mathcal{L}_{\text{ELBO}}$: